Need explanation and solution on how to solve problem.
Question 5 Exhibit 2-2 Given the following set of numbers: X1 = 2, X2 = 4, X3 = 6, X4-10 Refer to Exhibit 2-2. What is the value of ( Σ X) 2? a. 44 b. 156 O c. 484 Od. 480 Moving to another question will save this response.
Answer #1

〉 X=X1 + X2+X3 + X4 = 2+ 4 + 6 + 10 = 22

So,

)2 222 484

So,

Correct option:

c. 484